OpenStudy (anonymous):

What polynomial identity should be used to prove that 21=25-4? A: Difference of cubes B: Difference of Squares C: Square of Binomial D: Sum of Cubes

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

b) i think but im not sure let joanne answer

OpenStudy (joannablackwelder):

Hmm, 25-4 looks like difference of squares, but not sure how that helps you prove it.

OpenStudy (anonymous):

I thought it was c but im not positive.

OpenStudy (zpupster):

i had to look this one up. a 2 - b 2 = (a+b)(a-b) (Difference of squares) 5^2 - 2^2 lets see (5+2)(5-2) = ?

OpenStudy (joannablackwelder):

Yeah, @zpupster , that works. I'm just not sure how it proves the statement.